With the completion of the determinations of % potassium, % iron, and % oxalate in the crystals, you may calculate the % water. The percentage compositionof the crystals, KxFe(C2O4)y · zH2O, has then been completely determined experimentally. The simplest formula (x,y,z) can now be calculated from the the percentage composition. Once the formula is know it is then possible to calculate the percent yield of product that was obtained in the preparation and purification of the crystals.

From Part A:

Mass of KxFe(C2O4)y · zH2O prepared : 4.800 g
Mass of FeCl3 : 1.60 g

From Part B:

% Potassium in compound : 19.43 %
% Iron (from ion exchange & titration vs. NaOH) : 10.95 %

From Part C:

% Oxlate : 43.45 %

Using your chemical knowledge and literature references (dont forget to include the references in your lab report and discusUnsure about my first two answers so if you could check those it would be greatly appreciated!
Answer #1

Q1. % water of hydration = 100 - (% Potassium + % Iron + % Oxalate)

                                 = 100 - (19.43 + 10.95 + 43.45)

                                = 26.17 %

Assuming 100 g of sample is taken:

Elememt/molecule Mass (g) in 100 g of compound molar mass (in g/mol) mol present (=mass/molar mass) relative ratio of mole relative ratio in whole number
Potassium (K) 19.43 39.10 0.4969 (0.4969/0.1961) = 2.52 3 = x
Iron (Fe) 10.95 55.84 0.1961 (0.1961/0.1961) = 1.00 1
Oxalate (C2O4)2- 43.45 88.02 0.4936 (0.4936/0.1961) =2.52 3 = y
Water (H2O) 26.17 18.01 1.453 (1.453/0.1961) = 7.41 7 = z

So, the empirical formula : K3Fe(C2O4)3 · 7H2O

Q2. Mass of FeCl3 : 1.60 g = (1.60/162.2) = 0.00986 = mole of Fe

(As, molar mass of FeCl3 = 162.2 g/mol).

For Q3. And Q4. Molar mass of K3Fe(C2O4)3 · 7H2O= 3 x 39.10 + 55.84 + 3 x88.01 + 7 x 18.01

                                          = 563.11 g/mol   

Q 3. 1 mole of FeCl3 contain 1 mole Fe , And

as 1 mole of Fe is present in the 1 mole of complex,

Therefore, 0.00986 mole of Fe should give 0.00986 mole of complex = theoretical moles

Q 4. Actual mole synthesized = (4.800 g) /(563.11 g/mol) = 0.00852

Q 5. Percent yield = (actual moles) /theoretical moles

                              = (0.00852)/ (0.00986) = 0.8645

So, in % = 86.45